# LIA: Privacy-Preserving Data Quality Evaluation in Federated Learning Using a Lazy Influence Approximation

### v3

@@ 4.2 Filtration Metrics: Recall, Precision, Accuracy
4.3 Privacy
As expected, there is a trade-off between privacy and filtration quality (see Figure 4, vertical axis, where $\varepsilon$ refers to the privacy guarantee for both the training and validation data/participant votes). Nevertheless, Figure 4 demonstrates that our approach can provide reliable filtration, even under really strict, worst-case privacy requirements ($\varepsilon=1$, which is the recommended value in the DP literature [54], $\delta=10^{-5}$). Importantly, our decentralized framework allows each participant to compute and tune his own worst-case privacy guarantee a priori (see Section 3.2.3).
@@ 4.4 Robustness
We evaluated our approach for varying mislabeling percentage (0%, 10%, 20%, 30%, 40%), varying levels of non-IID-ness ($\alpha=0.1,1,10$), and a setting where not all data in a mislabeled batch are corrupt (90% corrupt). In all cases, the proposed approach enabled effective filtering (recall of $>90\%$). Please see the Appendix for detailed results.
5 Conclusion
@@
Privacy protection is a core element of Federated Learning. However, this privacy also means that it is significantly more difficult to ensure that the training data actually improves the model. Mislabeled, corrupted, or even malicious data can result in a strong degradation of the performance of the model (Figure 1) and privacy protection makes it significantly more challenging to identify the cause. We propose a practical influence approximation (‘lazy influence’) to be used as a ‘rite of passage’, allowing for effective filtering (recall of $>90\%$, and even up to 100%), while providing strict, worst-case $\varepsilon$-differential privacy guarantees ($\varepsilon\leq 1$) for both the training and validation data.
Appendix A Appendix / Supplemental Material
Contents
This appendix covers further details of our work, which have been omitted due to space limitations. Specifically:
  1. 1.
    In Section B we describe the methodology of our benchmarks.
  2. 2.
    In Section C we discuss the datasets used, the selected hyper-parameters, and other implementation details.
  3. 3.
    In Section D we describe the potential positive and negative societal impact of our work.
  4. 4.
    In Section E we briefly discuss the limitations of our work.
  5. 5.
    In Sections F we provide additional discussion of our simulation results.
Appendix B Setting
We consider a classification problem from some input space $\mathcal{X}$ (e.g., features, images, etc.) to an output space $\mathcal{Y}$ (e.g., labels). In a Federated Learning setting, there is a center $C$ that wants to learn a model $M(\theta)$ parameterized by $\theta\in\Theta$, with a non-negative loss function $L(z,\theta)$ on a sample $z=(\bar{x},y)\in\mathcal{X}\times\mathcal{Y}$. Let $R(Z,\theta)=\frac{1}{n}\sum_{i=1}^{n}L(z_{i},\theta)$ denote the empirical risk, given a set of data $Z=\{z_{i}\}_{i=1}^{n}$. We assume that the empirical risk is differentiable in $\theta$.The training data are supplied by a set of data holders.
(a) $\bm{\alpha}\rightarrow\infty$
(b) $\bm{\alpha}\rightarrow 100$
(c) $\bm{\alpha}\rightarrow 10$
(d) $\bm{\alpha}\rightarrow 0.1$
(e) $\bm{\alpha}\rightarrow 0.01$
(f) $\bm{\alpha}\rightarrow 0$
Figure 5: Dirichlet distribution visualisation for 10 classes, parametrized by $\bm{\alpha}$. $\bm{\alpha}$ controls the concentration of different classes. Each row represents a participant, each color a different class, and each colored segment the amount of data the participant has from each class. For $\bm{\alpha}\rightarrow\infty$, each participant has the same amount of data from each class (IID distribution). For $\bm{\alpha}\rightarrow 0$, each participant only holds data from one class. In this work, we use $\bm{\alpha}\rightarrow 0.1$ for a non-IID distribution.
B.1 Non-IID Setting
The main hurdle for Federated Learning is that not all data is IID. Heterogeneous data distributions are all but uncommon in the real world. To simulate a Non-IID distribution, we used Dirichlet distribution to split the training dataset as in related literature [28, 40, 27, 68]. This distribution is parameterized by $\alpha$, which controls the concentration of different classes, as visualized in Figure 5. This work uses $\alpha\rightarrow 0.1$ for a non-IID distribution, as in related literature (e.g., [68]).
B.2 Exact Influence
In simple terms, influence measures the marginal contribution of a data point on a model’s accuracy. A positive influence value indicates that a data point improves model accuracy, and vice-versa. More specifically, let $Z=\{z_{i}\}_{i=1}^{n}$, $Z_{+j}=Z\cup z_{j}$ where $z_{j}\not\in Z$, and let
$\hat{R}=\min_{\theta}R(Z,\theta)\text{\quad and\quad}\hat{R}_{+j}=\min_{\theta}R(Z_{+j},\theta)$
where $\hat{R}$ and $\hat{R}_{+j}$ denote the minimum empirical risk of their respective set of data. The influence of datapoint $z_{j}$ on $Z$ is defined as:
$\mathcal{I}(z_{j},Z)\triangleq\hat{R}-\hat{R}_{+j}$ (4)
Despite being highly informative, influence functions have not achieved widespread use in Federated Learning (or Machine Learning in general). This is mainly due to the computational cost. Equation 4 requires complete retraining of the model, which is time-consuming, and very costly; especially for state-of-the-art, large ML models. Moreover, specifically in our setting, we do not have direct access to the training data. In the following section, we will introduce a practical approximation of the influence, applicable in Federated Learning scenarios.
B.3 Influence Approximation
The first-order Taylor approximation of influence, adopted by [35] (based on [12]), to understand the effects of training points on the predictions of a centralized ML model. To the best of our knowledge, this is the current state-of-the-art approach to utilizing the influence function in ML. Thus, it is worth taking the time to understand the challenges that arise if we adopt this approximation in the Federated Learning setting.
Let $\hat{\theta}=\arg\min_{\theta}R(Z,\theta)$ denote the empirical risk minimizer. The approximate influence of a training point $z_{j}$ on the validation point $z_{val}$ can be computed without having to re-train the model, according to the following equation:
$\mathcal{I}_{appr}(z_{j},z_{val})\triangleq-\nabla_{\theta}L(z_{val},\hat{\theta})H^{-1}_{\hat{\theta}}\nabla_{\theta}L(z_{j},\hat{\theta})$ (5)
where $H^{-1}_{\hat{\theta}}$ is the inverse Hessian computed on all the model’s training data. The advantage of Equation 5 is that we can answer counterfactuals on the effects of up/down-scaling a training point, without having to re-train the model. One can potentially average over the validation points of a participant, and/or across the training points in a batch of a contributor, to get the total influence.
B.4 Challenges
(a) Random Label IID
(b) Label Shift Non-IID ($\alpha=0.01$)
Figure 6: Model accuracy relative to different mislabel rates (5% - 50%). These models have been trained over 25 communication rounds and 100 participants. We compare a centralized model with no filtering of mislabeled data (blue) to an FL model under perfect (oracle) filtering (orange). Note that the lower accuracy on the Non-IID setting is due to the fact that we are considering the most extreme non-IID case. This is where the majority of the participants have access to at most 1 class.
Consider Figure 6 as a motivating example. In this scenario, we have participants with corrupted data. Even a very robust model (ViT) loses performance when corruption is involved. This can also be observed in the work of [37]. Filtering those corrupted participants (orange line) restores the model’s performance.
While Equation 5 can be an effective tool in understanding centralized machine learning systems, it is ill-matched for Federated Learning models, for several key reasons.
To begin with, evaluating Equation 5 requires forming and inverting the Hessian of the empirical risk. With $n$ training points and $\theta\in\mathbb{R}^{m}$, this requires $O(nm^{2}+m^{3})$ operations [35], which is impractical for modern-day deep neural networks with millions of parameters. To overcome these challenges, [35] used implicit Hessian-vector products (HVPs) to more efficiently approximate $\nabla_{\theta}L(z_{val},\hat{\theta})H^{-1}_{\hat{\theta}}$, which typically requires $O(p)$ [35]. While this is a somewhat more efficient computation, it is communication-intensive, as it requires transferring all of the (either training or validation) data at each FL round. Most importantly, it can not provide any privacy to the users’ data, an important, inherent requirement/constraint in FL.
Finally, to compute Equation 5, the loss function has to be strictly convex and twice differentiable (which is not always the case in modern ML applications). A proposed solution is to swap out non-differentiable components for smoothed approximations [35], but there is no quality guarantee of the influence calculated in this way.
Appendix C Implementation details
This section describes the base model used in our simulations and all hyper-parameters. Specifically, we used a Visual Image Transformer (VIT) [16, 61]. The basis of our model represents a model pre-trained on ImageNet-21k at 224x224 resolution and then fine-tuned on ImageNet 2012 at 224x224 resolution. All hyper-parameters added or changed from the default VIT hyper-parameters are listed in Table 1 with their default values. The following hyper-parameters have been added to support our evaluation technique:
  • Random Factor: this coefficient represents the amount of corrupted data inside a corrupted batch.
  • Final Evaluation Size: an a priori separated batch of test data to evaluate model performance.
  • Parameters to Change: number of parameters (and biases) in the last layer of the model.
For the HAR dataset we have used a simple two-layer fully connected neural network. This network has not been pretrained like the previous one. With this network we chose to omit regularization and the detailed hyper-parameters are listed in Table 1. [67]
Regarding reproducibility, we ran the provided (in the supplementary material) code for each dataset with seeds from the range of $0-7$.
CIFAR10 CIFAR100 HAR
No. of Participants 100 100 500
Batch Size 100 250 2000
Validation Size 50 50 500
Random Factor 0.9 0.9 1.0
Warm-up Size 600 4000 7352
Final Evaluation Size 2000 2000 2947
Load Best Model False False False
Parameters to Change 7690 76900 36358
Learning Rate 0.001 0.001 0.001
Train Epochs 3 3 20
Weight Decay 0.01 0.01 None
Table 1: Table of hyper-parameters.
C.1 Termination Condition
Different termination conditions have been used for our proposed solution and to retrain the exact influence. Our solution has only one termination condition, that is the number of local epochs $k$.
C.2 Computational Resources
All simulations were run on two different systems:
  1. 1.
    Intel Xeon E5-2680 – 12 cores, 24 threads, 2.5 GHz – with 256 GB of RAM, Titan X GPU (Pascal)
  2. 2.
    Ryzen 5900X - 12 cores, 24 threads, 3.7 GHz - with 32 GB of RAM, RTX 2080 Super
Appendix D Societal Impact
Privacy advocacy movements have, in recent years, raised their voices about the potential abuse of these systems. Additionally, legal institutions have also recognized the importance of privacy, and have passed regulations in accordance, for example, the General Data Protection Regulation (GDPR). Our work provides practical privacy guarantees to protect all parties, with minimal compromise on performance. Furthermore, we allow data holders and collectors to be paid for their contribution in a joint model, instead of simply taking the data. Such incentives could potentially help speed up the growth of underdeveloped countries, and provide more high-quality data to researchers (as an example application, consider paying low-income farmers for gathering data in crop disease prevention [47]).
Appendix E Limitations
The main limitation of our approach is that if the optimizer does not produce a good enough gradient, we cannot get a good approximation of the direction the model is headed for. The result of this is a lower score, and therefore a potentially inaccurate prediction.
(a) Label shift Non-IID ($\alpha=0.01$)
(b) Random Label IID
Figure 7: Model accuracy over 25 communication rounds with a 30% mislabel rate on CIFAR-10. We compare a centralized model with no filtering (blue) to an FL model under perfect (oracle) filtering (orange), KRUM (red), Trimmed-mean (purple), and our approach (green). Note that the jagged line for KRUM is because only a single gradient is selected instead of performing FedAvg.
Another potential limitation is the filtering of “good” data. These data may be correctly labeled, but including it does not essentially provide any benefit to the model, as can be shown by the accuracy scores in Figure 7. While this allows us to train models of equal performance with a fraction of the data, some participants may be filtered out, even though they contribute accurate data. This might deter users from participating in the future.
Corruption
0% 10% 20% 30% 40%
Accuracy 85.2 +- 2.1 % 86.4 +- 1.2 % 87.2 +- 1.3 % 87.4 +- 1.1 % 89.0 +- 1.2 %
Recall - 97.0 +- 0.5 % 96.4 +- 0.7% 98.1 +- 0.6% 97.5 +- 0.7 %
Table 2: Filtering performance on varying percentages of corrupt participants on the Human Activity Recognition dataset for highly non-IID. Following the same strict worst-case differential privacy guarantees ($\varepsilon\leq 1,\delta=10^{-5}$).
Figure 8: Effect of batch size to the percentage of the times the sign of the proposed Lazy Influence Approximation (LIA) matches the sign of the exact influence, for varying differential privacy guarantees ($\varepsilon$ in the $x$-axis). Comparing 8(a) to the other two figures, we can observe that there is a certain threshold of data that needs to be passed for lazy influence to be effective. After this threshold has been reached, adding more data only gives marginal improvement as can be seen by comparing 8(b) and 8(c).
Appendix F Numerical Results
Filtration Metrics
Distribution Recall Precision Accuracy
CIFAR 10 IID 97.08 ± 3.51 % 91.91 ± 7.15 % 96.38 ± 2.83 %
Non-IID 93.75 ± 5.12 % 69.02 ± 6.28 % 85.00 ± 3.28 %
CIFAR 100 IID 99.17 ± 2.20 % 97.96 ± 2.30 % 99.12 ± 1.27 %
Non-IID 92.50 ± 5.71 % 55.41 ± 3.94 % 75.12 ± 3.76 %
HAR IID 100.00 ± 0.0% 100.00 ± 0.0% 100.00 ± 0.0%
Non-IID 95.77 ± 2.1% 71.71 ± 1.3% 87.40 ± 1.1%
Table 3: Filtering performance on various datasets, including real-data on Human Activity Recognition, for IID and highly non-IID setting ($\alpha\rightarrow 0.1$, i.e., 3 classes per participant for HAR). 100 participants, 30% mislabeling rate. Strict worst-case differential privacy guarantees ($\varepsilon\leq 1,\delta=10^{-5}$).
We provide detailed results that include both the means and standard deviations. The metrics can be found in Table 3. The following subsections provide a more comprehensive analysis of the results, summarized in the main text due to space limitations.
We have conducted initial testing on CIFAR10 and CIFAR100 to explore the impact of various parameters on our model’s performance. Our results, illustrated in Figures, 10, and 11 demonstrate the effect of both learning rate and number of epochs on filtration performance.
We observe a balance between recall and accuracy that varies based on the parameters. This balance can be seen in both the CIFAR10 and CIFAR100 datasets. Additionally, the best parameters for IID and Non-IID may differ. For instance, the best recall for Non-IID and IID is achieved with different parameter pairs, and CIFAR100 also has a distinct parameter pair for IID compared to Non-IID.
Finally, we examine the impact of various privacy guarantees ($\varepsilon$) and larger problem dimensions in Figure 9. Our findings show that a smaller federation is needed to achieve the same level of performance when data is IID, compared to when it is Non-IID.
(a) Recall
(b) Recall
(c) Precision
(d) Precision
(e) Accuracy
(f) Accuracy
Figure 9: Recall (top), Precision (middle), and Accuracy(Bottom) on CIFAR 10, non-IID (left), IID (right), for increasing problem size (number of participants), and varying privacy guarantees ($\varepsilon$ – lower $\varepsilon$ provides stronger privacy).
(a) Recall
(b) Recall
(c) Precision
(d) Precision
(e) Accuracy
(f) Accuracy
Figure 10: Recall (top), Precision (middle), and Accuracy(Bottom) on CIFAR 100 (left) and CIFAR 10 (right), IID, for different parameter pairs of learning rate and epoch count.
(a) Recall
(b) Recall
(c) Precision
(d) Precision
(e) Accuracy
(f) Accuracy
Figure 11: Recall (top), Precision (middle), and Accuracy(Bottom) on CIFAR 100 (left), and CIFAR 10 (right), non-IID, for different parameter pairs of learning rate and epoch count.

### v4

@@ IV-A Model Performance
IV-B Filtration Metrics: Recall, Precision, Accuracy
IV-C Privacy
As expected, there is a trade-off between privacy and filtration quality (see Figure 4, vertical axis, where $\varepsilon$ refers to the privacy guarantee for both the training and validation data/participant votes). Nevertheless, Figure 4 demonstrates that our approach can provide reliable filtration, even under really strict, worst-case privacy requirements ($\varepsilon=1$, which is the recommended value in the DP literature [54], $\delta=10^{-5}$). Importantly, our decentralized framework allows each participant to compute and tune his own worst-case privacy guarantee a priori (see Section III-B3).
@@ IV-D Robustness
We evaluated our approach for varying mislabeling percentage (0%, 10%, 20%, 30%, 40%), varying levels of non-IID-ness ($\alpha=0.1,1,10$), and a setting where not all data in a mislabeled batch are corrupt (90% corrupt). In all cases, the proposed approach enabled effective filtering (recall of $>90\%$). Please see the supplement VII-A for detailed results.
IV-E Minority Distributions
An important concern when filtering data is to ensure that clients with minority distributions will not be filtered out. Our results thus far demonstrate that our proposed approach achieves high quality filtering (outperforming established baselines), even on very hard non-IID settings (as low as just 3 classes per participant). Yet, these are aggregate results. To investigate specifically the case where clients have data from minority distributions, we consider the most extreme scenario, where only one client has data from a specific class (consider e.g., a hospital that treats a rare disease).
Specifically, we consider a scenario where every validator has data from all but one classes (leave-one-out, i.e., $\mathcal{Y}\setminus\{y_{out}\}$), drawn from a Dirichlet distribution, in a highly non-IID setting (as in our other simulations). Let a new incoming client $A$ have data from all class (including $y_{out}$). Let 5%, 10%, 15%, 20% of client $A$’s data belong to his unique class $y_{out}$, while the rest are drawn from classes $\mathcal{Y}\setminus\{y_{out}\}$ using the Dirichlet distribution. Figure 5 shows that, with high probability ($0.63$ for 5%, $0.54$ for 20%), client $A$ will not be filtered out. Our intuition is that client $A$ also has data that (partially) overlap with the validators (even in highly non-IID settings), and this overlap results in positive influence, in spite of their minority class data.
Figure 5: Probability that a client with data from a unique distribution (see Section IV-E) will be accepted in the federation. With high probability ($>50\%$) the client will not be filtered out, even in the extreme case of having 20% of data from a unique distribution. Results averaged over 512 runs.
V Conclusion
@@
The quality of a model obtained through Federated Learning can only be as good as the quality of the contributed data. Mislabeled, corrupted, or even malicious data can result in a strong degradation of the performance of the model (Figure 1). However, it is challenging to ensure this quality while preserving the privacy guarantees that are a hallmark of federated learning, and at the same time ensure that data from minority classes is not discarded. We propose a practical influence approximation (‘lazy influence’) to be used as a ‘rite of passage’, allowing for effective filtering (recall of $>90\%$, and even up to 100%), while providing strict, worst-case $\varepsilon$-differential privacy guarantees ($\varepsilon\leq 1$) for both the training and validation data.